### v1

# Towards Open-World Retrieval-Augmented Generation on Knowledge Graph: A Multi-Agent Collaboration Framework

Abstract
Large Language Models (LLMs) have demonstrated strong capabilities in language understanding and reasoning. However, their dependence on static training corpora makes them prone to factual errors and knowledge gaps. Retrieval-Augmented Generation (RAG) addresses this limitation by incorporating external knowledge sources, especially structured Knowledge Graphs (KGs), which provide explicit semantics and efficient retrieval. Existing KG-based RAG approaches, however, generally assume that anchor entities are accessible to initiate graph traversal, which limits their robustness in open world settings where accurate linking between the query and the entity is unreliable. To overcome this limitation, we propose AnchorRAG, a novel multi-agent collaboration framework for open-world RAG without the predefined anchor entities. Specifically, a predictor agent dynamically identifies candidate anchor entities by aligning user query terms with KG nodes and initializes independent retriever agents to conduct parallel multi-hop explorations from each candidate. Then a supervisor agent formulates the iterative retrieval strategy for these retriever agents and synthesizes the resulting knowledge paths to generate the final answer. This multi-agent collaboration framework improves retrieval robustness and mitigates the impact of ambiguous or erroneous anchors. Extensive experiments on four public benchmarks demonstrate that AnchorRAG significantly outperforms existing baselines and establishes new state-of-the-art results on the real-world question answering tasks.
1 Introduction
Large Language Models (LLMs) (Achiam et al. 2023; Yang et al. 2025; Liu et al. 2024) are typically defined as deep learning models with a massive number of parameters, trained on large-scale corpora in a self-supervised manner. Their internal parameterization allows for an implicit representation of external knowledge. During the inference stage, the Chain-of-Thought (CoT) (Wei et al. 2022; Trivedi et al. 2022; Zhou et al. 2022) method guides the models to “think step-by-step” through carefully designed prompts. It helps the models better handle logically complex or multi-step reasoning tasks, achieving a better performance in natural language processing tasks such as question answering (Gu et al. 2021). Despite these advances, LLMs are fundamentally limited by the incompleteness and inaccuracies in their static training corpora. This often results in factual hallucinations (Huang et al. 2025), especially in knowledge-intensive tasks, significantly hindering their reliability and deployment in real-world applications.
Figure 1: Illustration of the difference between the closed-world and the open-world setting in RAG.
To address these limitations, recent studies (Jeong et al. 2024; Xia et al. 2025) have investigated augmenting LLMs with external knowledge sources to improve factual consistency and reasoning. While fine-tuning (Hu et al. 2022; Li and Liang 2021; Li et al. 2023b), LLM with the external knowledge is computationally expensive and inflexible, a promising solution is Retrieval-Augmented Generation (RAG) (Lewis et al. 2020; Jiang et al. 2023b; Sarthi et al. 2024), which combines real-time retrieval from external sources with generative modeling. Among exiting methods, RAG on the Knowledge Graph (KG) (Jiang et al. 2023a; Ma et al. 2025; He et al. 2024; Chen et al. 2024; Guo et al. 2024) has become a research hotspot in the field, due to the fact that KGs can provide structured knowledge and explicit retrieval paths which benefit to enhancing the reasoning capabilities of LLMs. Typically, these KG-based RAG methods (Wang et al. 2025; Sun et al. 2023; Ma et al. 2024) can be divided in two steps: (1) identifying a query-aware anchor entity to initiate retrieval; (2) performing iterative retrieval to construct the reasoning paths. Through above steps, these methods can effectively exploit the external knowledge related to the query and improve the performance in the question answering task.
However, existing KG-based RAG methods almost follow the closed-world setting where the anchor entity is accessible and exists in the given KG. As shown in the Fig. 1 (a), the anchor entity “Mitt Romney” is pre-defined for the question “What university did Romney graduated from?”. The RAG methods can directly locate the target entity for the following retrieval. Actually, user questions usually conform to the open-world setting where the anchor entity is unavailable. In that case, existing methods extract the keywords in the query and identify the candidate anchor entity by semantic matching. Due to name abbreviations and aliases, these methods usually suffer from the issue caused by imprecise or partial matching as shown in Fig. 1(b). Then the challenge of the open-world RAG lies in how to accurately identify anchor entities to retrieve the relevant information.
To tackle this issue, we propose AnchorRAG, a multi-agent collaborative (MAC) framework that can perform RAG without the predefined anchor entities, thereby enabling more effective knowledge retrieval to enhance the LLM reasoning performance. Specifically, our framework follows a pipeline-based MAC framework, and a predictor agent first extracts the keywords from the given question and applies a semantic match model to obtain the candidate entities by their name description and structure neighborhood. Then, multiple retriever agents initiate parallel graph traversal with each candidate, conducting iterative retrieval to obtain the pivotal knowledge for the following reasoning. Finally, a supervisor agent will synthesize the retrieved evidence and determine whether an answer can be generated and formulate the retrieval process. The main contributions of this paper are summarized as follows:
  • General Aspect. We emphasize the importance of identifying the accurate anchor entity for the open-world RAG. By integrating multi-agent collaboration into the workflow of question answering, we can enhance the generality of RAG methods and improve the performance in the open-world setting.
  • Methodologies. For anchor identifying, we design an entity grounding strategy to locate the candidate anchors by entity name description and structural neighborhood. For knowledge retrieval, we propose a novel retrieval method with rough pruning and fine filtering, which can capture the resultful knowledge paths to boost the LLM reasoning.
  • Experimental Findings. Extensive experiments demonstrate that AnchorRAG consistently outperforms existing baselines on four public QA datasets, establishing state-of-the-art results on the real-world question answering tasks.
Figure 2: An overview of the AnchorRAG framework, including the collaborative workflows of multi-agent exploration for question answering.
2 Related Work
While advanced prompting methods like Chain-of-Thought (Wei et al. 2022) enhance LLM reasoning, their reliance on static internal knowledge can lead to factual errors and hallucinations that accumulate over multi-step inference. Retrieval-Augmented Generation (RAG) (Lewis et al. 2020; Semnani et al. 2023; He, Zhang, and Roth 2022) mitigates this by leveraging external knowledge, significantly improving performance on knowledge-intensive tasks like Knowledge Graph Question Answering (KGQA)  (Yih et al. 2016), where the knowledge graphs (KGs) (Bollacker et al. 2008; Auer et al. 2007; Vrandečić and Krötzsch 2014) provide large-scale structured facts and rich semantic relationships. These KG-based RAG methods can be divided into two main groups: KG-augmented Fine-tuning and KG-augmented In-context Learning.
2.1 KG-augmented Fine-tuning
KG-augmented fine-tuning directly integrates structured knowledge into open-source LLMs during training, operating at the entity, path, or subgraph levels (Zhang et al. 2025; Pan et al. 2024). At the entity level, methods like ChatDoctor (Li et al. 2023a) and PMC-LLaMA (Wu et al. 2024) employ instruction tuning (Wang et al. 2022) to enhance domain-specific knowledge understanding before fine-tuning on downstream tasks that contain vast entities and text descriptions. For path-level fine-tuning, RoG (Luo et al. 2023b) trains the LLM to generate valid relation paths, thereby aligning the reasoning process with the graph’s actual structure. The sources of the knowledge paths is diverse, ranging from the direct routes between question-related entities to complex logical rules extracted by graph retrieval models (Mavromatis and Karypis 2024) or heuristic strategies (Tan et al. 2024). To capture more comprehensive structural information, subgraph-level methods fine-tune LLMs on the entire local neighborhood, compressing them into embeddings (He et al. 2024) or transforming them into text sequences (Luo et al. 2023a).
2.2 KG-augmented In-context Learning
For state-of-the-art closed-source LLMs where fine-tuning is infeasible (Zhang et al. 2025), KG-augmented In-context Learning (ICL) is the primary strategy (Huang et al. 2023). Early methods like ToG (Sun et al. 2023) treat the LLM as an agent that executes sequential reasoning along the retrieved relational path. Given that this methods are prone to error accumulation and inefficient search, subsequent research has developed more advanced reasoning frameworks. To enhance reliability, Plan-on-Graph (Chen et al. 2024) introduced adaptive planning and self-correction mechanism, while Debate-on-Graph (Ma et al. 2025) proposed multi-path argumentation. Besides, recent methods, e.g., Fast Think-on-Graph  (Liang and Gu 2025), focused on optimizing search efficiency, they accelerates the reasoning process and improves accuracy by enabling LLMs to “think community-by-community” within the knowledge graph. To address knowledge graph incompleteness, ReKnoS (Wang et al. 2025) simplify reasoning by utilizing “super-relations”, while others, e.g., Generate-on-Graph (Xu et al. 2024), dynamically generate facts when external knowledge is unavailable. Nevertheless, almost of the existing methods are implemented for the closed-world question answering, which limits their performance in the real-world settings where accurate linking between the query and the entity is unreliable.
3 Method
3.1 Task Formulation
The fundamental unit of data in a knowledge graph $\mathcal{G}$ is the triples, i.e., $\mathcal{G}=\{(h,r,t)\mid h,t\in\mathcal{E},r\in\mathcal{R}\}$, where $\mathcal{E}$ and $\mathcal{R}$ respectively represent the set of entities and relations.The goal of Knowledge Graph Question Answering (KGQA) is to answer natural language questions by leveraging the structured triples within a knowledge graph. An unavoidable challenge for KGQA is how to design a synergy between the semantic understanding of LLMs and the structured facts of a knowledge graph, enabling both precise knowledge retrieval and powerful reasoning for the QA task.
3.2 Overview
As shown in the Fig. 2, our framework follows a pipeline-based MAC framework: anchor identifying by the predictor agent, knowledge searching by the retriever agents and question answering by the supervisor agent. Specifically, the proposed predictor agent first extracts the keywords from the given question and applies a semantic match model to obtain the candidate entities by their name description and structural neighborhood. Then, multiple retriever agents initiate parallel graph traversal for each candidate, performing iterative retrieval and obtaining the pivotal knowledge for the following reasoning. Furthermore, at each retrieval round, a supervisor agent aggregates the retrieved evidence and determines whether an answer can be generated and the retrieval process should be continued.
3.3 Anchor Entity Identifying
In our framework, the process of accurately identifying anchor entities from natural language questions is handled by the predictor agent. This process consists of the following two steps:
Candidate Entity Generation:
The predictor agent is associated with a LLM to recognize and extract keywords from the input question $q$, forming a keyword set $\mathcal{D}=\{d_{1},d_{2},\ldots,d_{n}\}$. For example, given the question $q$: ”What country bordering France contains an airport that serves Nijmegen?”, the predictor locates the keyword set $\mathcal{D}$, i.e., $\mathcal{D}=\{\text{France},\text{Nijmegen}\}$. In addition, considering there may exist certain spelling errors in the open-world questions, then the predictor will proofread the given question to filter out irrelevant distractors and correct potential errors, ensuring that the extracted keywords are clean and meaningful. Subsequently, it adopts a semantic match between these keywords and the entities set $\mathcal{E}$ of the external KG, generating an initial list of candidate anchors $\tilde{\mathcal{E}}$ through in-context learning. The prompt and in-context examples can be found in the Appendix.
Relation-Aware Entity Grounding:
The initial candidate set $\tilde{\mathcal{E}}$ often contains numerous entities with similar names that are irrelevant to the question’s context. Taking all of them into account will increase the overhead of subsequent retrieval and inevitably introduce noise. In this section, we argue that an entity is a more probable anchor if the semantic implied in its neighborhood is aligned with the given question. Therefore, the predictor agent presents a relation-aware grounding strategy to precisely identify the most relevant candidate entities. For each candidate entity $e\in\tilde{\mathcal{E}}$, the agent gathers its one-hop neighboring relations and obtains the set $\mathcal{R}(e)$. Then it invokes a pre-trained model SBERT (Reimers and Gurevych 2019) to encode the neighboring relations and the question, obtaining the relevance score $s(r_{i},q)$ between each relation $r_{i}\in\mathcal{R}(e)$ and the given question $q$ as follows:
$s(r_{i},q)=\left\langle\mathrm{SBERT}(r_{i}),\mathrm{SBERT}(q)\right\rangle,$ (1)
where $\left\langle\,\right\rangle$ denotes the inner product. Based on the above equation, we define the final relevance score of the candidate entity by aggregating the semantic information of its neighboring relations as follows:
$\mathrm{Score}(e,q)=\frac{1}{k}\sum_{r\in\mathcal{N}(e)}s(r,q),$ (2)
where $\mathcal{N}(e)$ denotes the set of top-$k$ relevant relations according to the Eq.(1). By scoring and ranking all candidates in the set $\tilde{\mathcal{E}}$, the predictor agent selects the top-$m$ entities with the higher relevance scores (e.g., France and Nijmegen) as the final anchor candidate set, which will be inputted to the following retriever agents for parallel searching.
3.4 Rough Pruning and Fine Filtering
Given the top-$m$ candidate anchor entities obtained from predictor agent, we assign each entity to an independent retriever agent. Thus there exist $m$ retrieval agents operating in parallel, each conducting iterative retrieval and reasoning based on its assigned anchor entity. At each iteration, these agents will retrieve the one-hop neighborhood of the given entity, where the neighborhood includes the neighboring relations and entities. Then a rough pruning and fine filtering method is applied to select more meaningful neighborhood for the question. Finally, the neighboring entities will be considered as the initial ones for the next retrieval iteration if necessary.
Neighboring Relations Pruning:
At the $t$-th retrieval iteration, each retriever agent explores candidate relations from its current given entities $e_{i}^{t}$. This agent retrieves all one-hop incoming and outgoing relations of the given entity from the knowledge graph $\mathcal{G}$. To reduce noise, the agent will filter out trivial or schema-level relations (e.g., type.object.name, common.*, sameAs), as well as previously visited relations. The remaining relations are scored by the LLM according to their relevance score with the given question. Specifically, each agent constructs a prompt containing the question, current anchor entity, and candidate relations, applying the LLM to assign the normalized scores. The prompt and in-context examples can be found in the Appendix. With the neighboring relations pruning, the top-$b$ relations are selected.
Neighboring Entities Pruning:
Given the selected relations, each agent retrieves the neighboring entities linked via these relations. Considering that each relation will have many neighboring entities, to bound the length of LLM input, we limit the number of entities per relation. Specifically, each candidate entity is ranked by the LLM with the prompt including the question, current anchor entity and the selected relation. The prompt and in-context examples can be found in the Appendix. Then top-$b$ entities are obtained for each relation, forming the knowledge triple set $\mathcal{T}(e_{i}^{t})$.
Candidate Triples Filtering:
To further identify the question-related facts implied in the knowledge graph, each agent applies a fine filtering method on the candidate triples $\mathcal{T}(e_{i}^{t})$, improving the performance in the real-world question answering. Specifically, the retrieval agent first integrates all candidate triples in $\mathcal{T}(e_{i}^{t})$ into a structured list and prompts a LLM to select the related triples according to the given question. The prompt and in-context examples can be found in the Appendix. Then the filtered triples are adopted as high-quality external knowledge which will be inputted to the following supervisor agent for answer generation. Furthermore, if the filtered triple set is empty, it means that the current exploration path can no longer yield related information for the given question. In that case, this agent will terminate its subsequent retrieval and provide no information for the supervisor at the current iteration. This “early-stopping” mechanism ensures the agent can adaptively waive unpromising knowledge paths, reducing the computational overhead for efficient retrieval-augmented generation.
   Dataset    WebQSP    GrailQA    CWQ    WebQuestions
   Hit@1    acc    Hit@1    acc    Hit@1    acc    Hit@1    acc
   Method    Qwen-Plus
   IO    63.3    42.2    33.2    26.6    33.2    33.2    56.3    43.2
   Chain-of-Thought    62.9    41.5    32.3    26.7    37.6    37.6    54.1    41.2
   Self-Consistency    63.0    41.2    33.8    28.4    39.2    39.2    52.7    40.1
   PoG    33.1    25.4    36.9    33.0    39.5    29.2    25.3    21.5
   ToG    66.1    46.3    41.9    35.3    39.8    39.8    56.1    43.7
   Ours    73.3    56.4    62.7    56.0    47.0    47.0    60.9    50.5
   Method    GPT-4o-mini
   IO    65.8    46.1    35.6    27.9    35.7    35.7    57.7    45.4
   Chain-of-Thought    62.5    41.6    31.0    25.8    34.5    34.5    53.6    41.2
   Self-Consistency    59.2    40.1    34.0    27.1    36.1    36.1    50.7    39.3
   PoG    55.3    36.4    39.0    34.2    36.1    36.1    41.8    33.3
   ToG    71.4    49.9    43.7    36.3    40.7    40.7    61.4    47.6
   Ours    74.1    57.4    63.4    56.8    44.8    44.8    60.0    50.0
Table 1: The results of different methods on various Datasets, using Qwen-Plus and GPT-4o-mini as the LLM. The best results are highlighted in bold. The suboptimal results are underlined.
3.5 Answer Generating
At the end of each iteration, all active retrieval agents input their resulting paths to a supervisor agent. The supervisor aggregates these paths and prompts the LLM to assess whether the current information is sufficient to answer the question. If so, it stops the reasoning process and returns the final answer. If not, only agents with valid paths are permitted to perform the next retrieval round. The prompt and in-context examples can be found in the Appendix. To avoid excessive computation overhead, the supervisor adopts the early termination if (1) no agents remain active, or (2) the maximum reasoning depth $L$ is reached. In such cases, the supervisor invokes LLM to directly answer the question by applying CoT reasoning with its internal knowledge.
4 Experiments
4.1 Dataset and Evaluation
In the experiments, we adopt four benchmark datasets to evaluate the performance of our framework, including three multi-hop Knowledge Graph Question Answering (Multi-Hop KGQA) datasets: WebQuestionSP (WebQSP) (Yih et al. 2016), GrailQA (Gu et al. 2021) and Complex WebQuestions (CWQ) (Talmor and Berant 2018), and one for Open-Domain Question Answering: WebQuestions (Berant et al. 2013). Freebase (Bollacker et al. 2008) is utilized as the external Knowledge Graph, which contains the complete knowledge facts that are necessary to answer questions in the above datasets. In addition, we apply the exact match (i.e., Hit@1) and accuracy (i.e., acc) as evaluation metrics for our method which are widely used in the field of question answering with RAG.
4.2 Implementation Settings
We extract all entities from Freebase, encode their name description into vectors using SBERT model (all-MiniLM-L6-v2, without fine-tuning) (Reimers and Gurevych 2019), and subsequently index them in a vector database (Douze et al. 2024) for downstream entity querying and matching. The knowledge graph derived from Freebase is deployed on a locally deployed Virtuoso (Erling and Mikhailov 2009) server for our experiments. We selected GPT-4o-mini and Qwen-Plus as the fundamental LLM in our framework. To prevent redundant exploration by agents and reduce computational overhead, we set both the maximum search depth $L$ and the expansion width $b$ to 3. This configuration aligns with the empirically optimal values identified in (Sun et al. 2023). Additionally, we fix the number of neighbor relations $k$ during anchor entity grounding to 5, and the number of retrieval agents $m$ is set to 3 for parallel multi-hop reasoning. The data and source can be found in appendix and will be available soon.
4.3 Baselines
Some widely used prompting strategies are selected as the baselines, including vanilla IO prompting (Brown et al. 2020), Chain-of-Thought (CoT) (Wei et al. 2022), and Self-Consistency (Wang et al. 2022). Moreover, we also compare the proposal with two representative RAG methods: Think-on-graph (Sun et al. 2023) and Plan-on-graph (Chen et al. 2024). It is worth noting that the above two methods are both implemented in the closed-world setting where the anchor entities are predefined for each question. Thus, to make a fair comparison, we apply the fundamental LLM to select the anchor entities instead of directly identifying them for the open-world question answering. Other settings and parameters remain unchanged.
4.4 Main Results
The main experimental results, as shown in Table 1, indicate that our proposed method achieves a significant performance advantage over all baseline methods when deployed with both the Qwen-Plus and GPT-4o-mini LLMs. Specifically, when applying Qwen-Plus, AnchorRAG achieves an average improvement by 10% and 11.2% on the metrics Hit@1 and Accuracy respectively, compared to the strongest baseline. Besides, with GPT-4o-mini model, our proposal also achieves the increases of 6.3% and 8.6%. We can also find that our method obtain an outstanding performance on the complex datasets like GrailQA, where its Hit@1 and Accuracy scores are increased substantially by 20.8% and 20.7%, respectively. The reason may be that the questions and the kG entities are semantically inconsistent in this dataset, and the anchor entities are more difficult to locate in this case. AnchorRAG can handle this issue by the multi-agent collaboration, thereby obtaining the state-of-the-art results. In addition, the performance of the RAG methods, e.g., ToG, consistently surpasses the methods solely relying on the internal knowledge, e.g., CoT. This demonstrates the effectiveness of the RAG methods in the question answering task, especially when dealing with the complex reasoning such as GrailQA and CWQ.
Dataset WebQSP GrailQA
normal open-world normal open-world
Method Qwen-Plus
[…]
Method GPT-4o-mini
[…]
Table 2: Performance comparisons (Hits@1) of different methods on the normal and open-world versions of WebQSP and GrailQA datasets, using Qwen-Plus and GPT-4o-mini as the backbone LLMs.
Method WebQSP GrailQA
AnchorRAG 74.1 63.4
w/o Entity Grounding 67.8 42.8
w/o Parallel Retrieval 72.8 60.2
w/o Triples Fine Filtering 71.6 62.8
Table 3: Ablation study of AnchorRAG: Effect of different components on the Hits@1 performance.
4.5 Robust Analyses
Given that existing public QA datasets primarily assume that the external KG contains all the necessary information for answering the given questions, which is not suitable for the open-world question answering task. Thus, to further evaluate the generality and robustness of our method, we constructed the open-world versions of the WebQSP and GrailQA datasets. Specifically, we first added the semantic noise by introducing some typos into the original questions to prevent exact matches with entities in the KG. Then we incorporated additional questions into the datasets whose answers could not be directly retrieved from the KG. The detailed construction process is provided in the appendix. The results on the self-constructed datasets are presented in Table 2, we can observe that our method shows strong robustness on this open-world scenario. For instance, on GrailQA, the performance of ToG using Qwen-Plus drops by 23.4%, whereas our method only drops by 11.6%. These results demonstrate that AnchorRAG can effectively address the open-world QA tasks and further enhance the generality of the RAG methods.
4.6 Ablation Studies
To validate the effectiveness of the core components within our proposed AnchorRAG framework, we conducted the following ablation studies on the WebQSP and GrailQA datasets, with the results presented in Table 3. Note that ‘w/o Entity Grounding’ removes the neighborhood information during anchor entity identification, ‘w/o Parallel Retrieval’ eliminates the multi-agent retrieval mechanism, replacing it with a single-path retrieval approach, and ‘w/o Triples Fine Filtering’ refers to the absence of fine semantic filtering for candidate triples during retrieval. The results show that each component can contribute to the question answering performance. Comparing with parallel retrieval and fine filtering, the entity grounding module has a greater impact on the performance, achieving the improvement by 6.3% and 20.6% on WebQSP and GrailQA respectively, which emphasizes the importance of precise anchor entities in the open-world question answering tasks.
Figure 3: The effect of the anchor identifying and knowledge retrieval on WebQSP.
4.7 Effect of the Retrieval
To further validate the effectiveness of our method in anchor identifying and knowledge retrieval, we compare the effect of retrieval of our method against two RAG baselines (PoG and ToG) and a variant (AnchorRAG w/o Parallel Retrieval). The results are shown in Fig. 3. The figures in the top row illustrate the accuracy of anchor identifying (marked in green), where our method significantly outperforms the baselines by grounding the entities with their textual and neighborhood information. The bottom row presents the proportion of the exact match for the answers, including the answers obtained by retrieval (marked in red) and those inferred by CoT reasoning (marked in orange). Notably, AnchorRAG achieves the highest retrieval rate at 49.6%, indicating that the multi-agent collaborative mechanism can effectively obtain the correct answer entities, which supports the following reasoning process to improve the QA performance.
4.8 Hyper-parameter Analyses
Setting $k=3$ $k=5$ $k=7$
$m=1$ 60.4 57.6 56.4
$m=2$ 63.6 63.3 61.0
$m=3$ 63.1 63.4 63.3
$m=4$ 62.6 63.1 62.4
Table 4: The results of AnchorRAG with different hyper-parameters using GPT-4o-mini on GrailQA.
We conducted the parameter sensitivity analyses on Grail-QA to investigate the optimal configuration for the number of neighborhood relations $k$ and retrieval agents $m$, where the results are shown in Table 4. We can find that retaining excessive neighboring relations $k$ for each candidate entity will lead to a performance loss, due to the fact that it may introduce noise and neglect the key structural information. Besides, a proper number of retriever agent $m$ can make AnchorRAG achieve the optimal results, which can reflect the effectiveness of the MAC framework. In addition, we can also find that more retrieval agents won’t always obtain the better results. Excessive agents will introduce uncertain retrieval information into the collaborative process, thereby impairing the performance.
4.9 Case Study
To intuitively demonstrate the superiority of our method in knowledge graph retrieval, we conduct a case study that compares our approach with the ToG baseline on a representative complex question from the CWQ dataset: “Who inspired F. Scott Fitzgerald, and who was the architect that designed The Mount?”. As shown in Fig. 4, we can observe that ToG can’t locate both the anchor entities, resulting in meaningless reasoning paths and generating an incorrect answer. In contrast, our multi-agent framework can effectively identify the anchor entities for the given questions by the entity grounding mechanism. With assigning multi retriever agents to independently explore different reasoning paths, AnchorRAG accurately locates the correct answer entity “Edith Wharton”. Additionally, this case highlights the robustness of our method. For example, although it initially considers an incorrect entity (“The Weapon”) as an anchor entity, it promptly detects the irrelevance of the following retrieval path with the rough pruning and fine filtering method, and terminates the corresponding agent, which prevents unnecessary computation overhead.
Figure 4: Comparisons of retrieval process between a baseline reasoning (ToG) and our method (AnchorRAG) on a complex question. The entities marked in yellow are anchor entities, marked in green are correct answer entities. The ones marked in blue and striped are invalid entities.
5 Conclusion
In this paper, we present AnchorRAG, a novel multi-agent collaborative RAG framework for open-world Knowledge Graph Question Answering. Unlike existing methods that rely on predefined anchor entities, AnchorRAG overcomes the bottleneck of unreliable entity linking by identifying and grounding plausible anchors without prior assumptions. Specifically, a predictor agent dynamically identifies candidate anchor entities by aligning the given questions with the candidate entities and initializes independent retriever agents to conduct parallel multi-hop explorations. Then a supervisor agent formulates the iterative retrieval strategy for these retriever agents and synthesizes the reasoning paths to generate the final answer. Extensive experiments on four public KGQA benchmarks demonstrate that AnchorRAG significantly outperforms the state-of-the-art baselines, validating the effectiveness of our MAC framework, especially when dealing with the complex multi-hop reasoning tasks.